OpenStudy (anonymous):

What is the volume of an oblique square pyramid with base edges 25ft and height 24ft?

OpenStudy (anonymous):

@ganeshie8

jhonyy9 (jhonyy9):

so can you draw it please ?

jhonyy9 (jhonyy9):

how is the oblique square ?

jhonyy9 (jhonyy9):

how many degree has the angle of deviation of this square ?

OpenStudy (anonymous):

volume of a a pyrmid is the same so just plug in those numbers.. \(\large V=\frac{1}{3}\cdot (area.of.base) \cdot (height) \) \(\large V=\frac{1}{3}\cdot (25^2) \cdot (24) \)

OpenStudy (anonymous):

 DEEZ NUTZ